shelby walked for awhile at 4 mph and then jogged the rest of the way at 8 mph. If she covered 56 miles in 10 hours how far did she walk, and how far did she jog?..any help would be GREAT! : )
Found 2 solutions by jasonlin00, rajagopalan:
Answer by jasonlin00(6)

She walked for 6 hours which was 24 miles.
Then she jogged for 4 hours which was 32 miles.
thats 56 miles in 10 hours.

Answer by rajagopalan(174)

Let Jogging hours = x
Walking hours = 10-x
Jogging distance = speeedxtime=8x
Walking distance=4(10-x)
Total distance=8x+4(10-x)=56
8x+4(10-x)=56
8x+40-4x=56
8x-4x=56-40
4x=16
x=4
Jog hours =4
Walk hours =6
Verify
Jog dist =4*8=32
Walk dist =6*4=24
Total = 56 miles
okay
good day